OpenStudy (anonymous):

The value of the 9 in 901, 532 is how many times the value of the 9 in 157, 900

OpenStudy (anonymous):

hundred

OpenStudy (radar):

thousand. 900, 9 * 10^2 900,000, 9*10^5 10^5/10^2 = 10^3= 1000